Engberg Company installs lawn sod in home yards. The company’s most recent monthly contribution format income statement follows: Amount Percent of Sales Sales $ 130,000 100 % Variable expenses 52,000 40 % Contribution margin 78,000 60 % Fixed expenses 16,000 Net operating income $ 62,000 Required: 1. What is the company’s degree of operating leverage? 2. Using the degree of operating leverage, estimate the impact on net operating income of a 29% increase in sales. 3. Construct a new contribution format income statement for the company assuming a 29% increase in sales.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Engberg Company
Answer 1 Amount $ Note
Sales       130,000.00 A
Less: Variable costs         52,000.00 B
Total Contribution margin         78,000.00 C=A-B
Less: Total fixed cost         16,000.00 D
Net operating income         62,000.00 E=C-D
Operating leverage                   1.26 F=C/E
Answer 2 Amount $
Increase in sales by 29% G
Impact in net operating income         22,620.00 H=E*F*G
Answer 3
Income statement Amount $
Sales       167,700.00 I= A*129%
Less: Variable cost         67,080.00 J=B*129%
Contribution margin       100,620.00 K=I-J
Less: Fixed cost         16,000.00 See E
Net operating income         84,620.00 L=K-E
Previous Net operating income         62,000.00 See E
Impact in net operating income         22,620.00 M=L-E
Hence proved.